OpenStudy (anonymous):

Find the area of the sector of the disk of radius r = 20 ft and with arc length s = 17 ft. Answer in units of ft^2

OpenStudy (anonymous):

any clue?

OpenStudy (anonymous):

\[\text{sectorArea}=\frac{17}{\pi 2\ 20}\pi 20^2\text{=}170 \text{ ft}{}^{\wedge}2\]

OpenStudy (anonymous):

sectorArea = (arc length/circle circumference)*(circle area)